# TDS-CLIP: Temporal Difference Side Network for Image-to-Video Transfer Learning

### v1

@@ Abstract
Video Swim-L(Liu et al. 2022b) IN-21K 32$\times$3$\times$4 0.6$\times$12 83.1 95.9
ViViT FE-L(Arnab et al. 2021) IN-21K 128$\times$3$\times$1 4.0$\times$3 81.7 93.8
MViT-B(Fan et al. 2021) Kinetics-600 32$\times$5$\times$1 0.2$\times$5 82.9 95.7
Uniformer V2-B/16(Li et al. 2023) CLIP-400M 8$\times$3$\times$4 0.2$\times$12 85.6 97.0
BIKE-L/14(Wu et al. 2023b) CLIP-400M 16$\times$3$\times$4 0.8$\times$12 88.1 97.9
ActionCLIP-B/16(Wang et al. 2023) CLIP-400M 32$\times$3$\times$10 0.6$\times$30 83.8 96.2
X-CLIP-B/16(Wang et al. 2023) CLIP-400M 16$\times$3$\times$4 0.6$\times$12 84.7 96.8
S-ViT-B/16(Zhao et al. 2023) CLIP-400M 16$\times$3$\times$4 0.7$\times$12 84.7 96.8
@@
EVL-B/16(Lin et al. 2022) CLIP-400M 8$\times$3$\times$1 0.4$\times$3 82.9 -
ST-Adapter-B/16(Pan et al. 2022) CLIP-400M 8$\times$3$\times$1 0.2$\times$3 82.0 95.7
ST-Adapter-B/16(Pan et al. 2022) CLIP-400M 32$\times$3$\times$1 0.6$\times$3 82.7 96.2
DUALPATH-B/16(Park et al. 2023) CLIP-400M 32$\times$3$\times$1 0.7$\times$3 85.4 97.1
AIM-B/16(Yang et al. 2023) CLIP-400M 8$\times$3$\times$1 0.2$\times$3 83.9 96.3
AIM-B/16(Yang et al. 2023) CLIP-400M 32$\times$3$\times$1 0.6$\times$3 84.7 96.7
DiST-B/16(Qing et al. 2023) CLIP-400M 8$\times$3$\times$1 0.2$\times$3 83.6 96.3
DiST-B/16(Qing et al. 2023) CLIP-400M 32$\times$3$\times$1 0.7$\times$3 85.0 97.0
M2-CLIP-B/16(Wang et al. 2024) CLIP-400M 32$\times$3$\times$4 0.8$\times$12 84.1 96.8
Side4Video-B/16(Yao et al. 2023) CLIP-400M 32$\times$3$\times$4 0.7$\times$12 84.2 96.5
@@
Table 2: Comparison of performance with state-of-the-art methods on Kinetics-400. “Cr.” and “Cl.” are the abbreviation for “spatial crops” and “temporal clips”.
TD-Adapter:
As shown in Fig. 3(b), the input feature of the TD-Adapter is denoted as $Z_{v}\in\mathbb{R}^{T\times N\times C}$. First, it is reshaped to a form $Z_{s}\in\mathbb{R}^{C\times T\times H\times W}$ suitable for 3D convolution. Then, a set of convolution operations is used for dimensionality reduction to obtain the feature $Z_{s}\in\mathbb{R}^{C/r\times T\times H\times W}$, $r$ denotes rate of reduction. Next, we capture the sensitive feature values of the local temporal dimension features using a 3D max-pooling operation, and a feature difference operation is employed to obtain dynamic features of local motion patterns.
$D_{z}=Z_{s}-Pool(Z_{s}),$ (4)
where $D_{z}$ denotes the feature difference. This method is commonly used in the early stages of convolutional models for efficient temporal modeling (Wang, She, and Smolic 2021; Li et al. 2020; Wang et al. 2021).
TDS-CLIP for video action recognition:
We integrate the proposed SME-Adapter and TD-Adapter into Side4Video(Yao et al. 2023) architecture, where the TD-Adapter, combined with a trainable ViT encoder, forms the TDS-CLIP block. We retain the CLS token shift(Liu et al. 2022a; Zhang, Hao, and Ngo 2021) operation to further improve the capacity for learning temporal information, offering the advantage of minimal additional memory overhead. The number of these blocks matches the frozen ViT encoders in the visual branch of CLIP and combined via side-connections. For action recognition tasks, we obtain the prediction scores by performing global average pooling on the output of the final TDS-CLIP block.
Experiments
Experimental Setup
To evaluate the effectiveness of our method, we assess it on several widely used action recognition benchmark datasets, including Something-Something V1 (SSv1) (Goyal et al. 2017), Something-Something V2 (SSv2) (Goyal et al. 2017), and Kinetics400 (Kay et al. 2017). SSv1 and SSv2 are commonly used datasets for evaluating the importance of temporal modeling. Kinetics400 is a comprehensive action recognition dataset spanning 400 different human actions. We use the CLIP pretrained ViT-B/16 and ViT-L/14 as spatial encoders. By adjusting the dimensions, our model balances memory usage and performance. Due to hardware dilemmas, unless otherwise specified, we use the corresponding ViT-B/16 or ViT-L/14 structure in the temporal encoder to form the TDS-CLIP framework, employing a sparse frame sampling strategy of 8 or 16 frames during both training and inference.
@@
(a) Study of component effectiveness.
@@
(b) The impact of local frame sample number.
$\alpha$ / $\beta$ Top-1(%) Acc.
0.5 58.4
1.0 59.4
1.5 58.9
(c) The impact of fusion rate.
Method Top-1(%) Acc.
Pool 59.4
Conv 59.3
Direct 57.3
(d) The impact of temporal difference module.
@@
(e) The impact of kernel size in pooling layers.
@@
(f) The impact of dimensionality.
Table 3: Ablation study results on Something-Something V1. Unless otherwise specified, all models use ViT-B/16 with 8 frames under the single view protocol.
Comparison with State-of-the-art
Results on Something-Something V1&V2:
Tab. 1 shows the comparison results with other state-of-the-art methods on SSv1 and SSv2. Both fully fine-tuning and frozen CLIP parameter methods, our approach achieves competitive results on both SSv1 and SSv2. For instance, in the fully fine-tuning approach, UniFormer-V2-L/14(Li et al. 2023) achieves the best classification results with the ViT-L/14 backbone, while our TDS-CLIP demonstrates superior performance with the same backbone (63.0% vs. 60.5% on SSv1; 73.4% vs. 72.1% on SSv2) while using fewer frames. In the frozen CLIP parameter methods, Side4Video-B/16(Yao et al. 2023) uses the same side network structure; in contrast, TDS-CLIP achieves better results with the same frame and sampling settings (61.0% vs. 60.7% on SSv1; 72.1% vs. 71.5% on SSv2), indicating that the proposed adapter enables effective temporal modelling. Additionally, it can be observed that the ViT-L/14 backbone outperforms the ViT-B/16 backbone under the same settings. These experimental results show that the proposed TD-Adapter and SME-Adapter are effective methods for enhancing temporal modeling and facilitating the transfer of knowledge from images to videos in frozen CLIP models.
Results on Kinetics-400:
To evaluate the generalization performance of our method, we conduct experiments on the K400 dataset. Compared to SSv1 and SSv2, the K400 emphasizes learning appearance features. As shown in Tab. 2, our method underperforms BIKE-L/14 (Wu et al. 2023b) in terms of appearance feature learning capability when compared to fully fine-tuned approaches. However, with 8-frame inputs, our performance surpasses that of ActionCLIP-B/16 (Wang et al. 2023), MViT-B (Fan et al. 2021), and Video Swin-L (Liu et al. 2022b) with 32-frame inputs, while also maintaining lower computational cost. When compared to methods that freeze CLIP, our performance with 32-frame inputs is lower than DiST-B/16 (Qing et al. 2023), AIM-B/16 (Yang et al. 2023), and DUALPATH-B/16 (Park et al. 2023), which possess independent spatial feature learning mechanisms. In contrast, TDS-CLIP focuses more on temporal feature learning and modeling. Notably, we achieve a competitive advantage when compared to methods like DiST-B/16 (Qing et al. 2023), AIM-B/16 (Yang et al. 2023), ST-Adapter-B/16 (Pan et al. 2022), and EVL-B/16 (Lin et al. 2022) with 8-frame inputs, indicating that TDS-CLIP captures more comprehensive video action information with fewer input frames. Lastly, we also evaluate the performance of TDS-CLIP with an 8-frame ViT-L framework, demonstrating that larger-scale frameworks are effective in improving performance. The experimental results demonstrate that TDS-CLIP not only excels in datasets with strong temporal correlations but also performs well in datasets with significant appearance background correlations. With these observations, we can conclude that TDS-CLIP has the dual advantage of spatial modeling and temporal modeling, showcasing robust generalization capabilities.
Ablation and Analysis
In this section, unless otherwise specifed, we use ViT-B/16 as the backbone and the 8 input frames on Something-Something V1.
Study of component effectiveness.
We explore the effectiveness of the proposed adapter and the results are shown in Tab. 3(a). The conventional 3D Conv is used to replace the TD-Adapter for the baseline of comparison. We can observe that the classification accuracy is improved by 1.0% after adding TD-Adapter, which proves that TD-Adapter can effectively improve the temporal modelling capability. Then, we add SME-Adapter, and the accuracy is improved even further, which implies that the SME-Adapter guides the side network to learn the motion features to further enhance the temporal modelling capability of the model, and proves that TD-Adapter and SME-Adapter are complementary to each other in terms of performance improvement.
The impact of local frame sample number.
We compare the performance impact of different numbers of local sampling video frames, and the results are shown in Tab. 3(b). 1 indicates that the SME-Adapter is not enabled, and serves as the baseline for comparison. The accuracy improves by 0.5% when the number is increased to 3 and by 1.4% when the number is increased to 5. This shows that increasing the number of local sampling frames can effectively improve the recognition performance. We do not test larger quantities because the total number of video frames used for testing is not large and bigger quantities may destroy the sparse sampling efficiency.
The impact of fusion rate.
The fusion ratio represents the proportion of the proposed method in feature fusion. The test results, as shown in Tab. 3(c), indicate that performance is optimal under equal-ratio fusion, while the performance deteriorates when the output feature ratio of the SME-Adapter is either too high or too low. This shows that the appropriate equal-ratio feature fusion method can well exploit the ability of SME-Adapter to learn and capture motion features.
Ablations for TD-Adapter.
We explore the impact of different core modules of the TD-Adapter on performance, including temporal convolutional layer enhancement(Wang et al. 2024) and direct feature difference calculation, as shown in Tab. 3(d). Using pooling layers achieves a classification accuracy of 59.4%, while temporal convolutional layer enhancement yields a slightly lower accuracy, decreasing by 0.2%. Direct feature difference calculation yields the lowest accuracy, as this approach tends to cause the loss of spatial features. It is noteworthy that pooling is a parameter-free operation, which helps reduce the number of parameters.
Furthermore, we test different kernel sizes for the pooling layers, specifically k=3, 5, and 7, as illustrated in Tab. 3(e). The classification accuracy decreases as the kernel size increases, indicating that an excessively large kernel size may disrupt the temporal relationships between features, therefore, larger kernel sizes should not be selected.
The impact of dimensionality.
In the transformer architecture, adjusting the dimensionality can effectively control the model’s memory consumption and impact its performance. Generally, increasing the dimensionality enhances the model’s complexity and memory usage, while also improving the modeling capacity of different layers. We follow the baseline comparison methods to test the impact of various dimensionality settings on performance, as shown in Tab. 3(f). The results show that the model achieves the best classification performance at 320 dimensions, while both excessively large or small dimensions constrain performance.
Layer 1-6 Layer 7-12 SME-Adapter Top-1(%) Acc.
58.8
58.6
59.1
59.4
Table 4: Study on the location of TD-Adapter layers.
Study on the location of TD-Adapter layers.
@@
We study the location of the TD-Adapter within the structure, with the results shown in Tab. 4. We experiment with adding the TD-Adapter to the first half of the network (layers 1-6) and the second half (layers 7-12). Replacement with normal 3D convolution when not using TD-Adapter. The table shows that while using more TD-Adapters generally leads to better classification results, adding them to shallower layers provides greater benefits than those obtained from deeper layers. Additionally, the performance is further improved when the SME-Adapter is also used.
@@
Figure 5: Visualisation of activation maps for features. Blue colour indicates larger magnitude of feature activation values and red colour indicates smaller magnitude of feature activation values.
Visualisation Related.
@@
As show in Fig. 5, we further visualize the activation maps of features in the side network with and without using the SME-Adapter. It is evident that the SME-Adapter is more sensitive to the motion areas of the target and does not show a bias towards background information. For example, our SME-Adapter focuses more on the motion region of the hand that interacts with the object, rather than all the targets such as flower pots, water bottles, or mobile phones that appear in the scene, which is more in line with the human perception. This indicates that the proposed SME-Adapter facilitates the learning of motion features in videos, enabling the model to focus more accurately on dynamically moving target objects.
Conclusion

### v2

@@ IV-A Experimental Setup
Video Swim-L [46] IN-21K 32$\times$3$\times$4 0.6$\times$12 83.1 95.9
ViViT FE-L [45] IN-21K 128$\times$3$\times$1 4.0$\times$3 81.7 93.8
MViT-B [47] Kinetics-600 32$\times$5$\times$1 0.2$\times$5 82.9 95.7
Uniformer V2-B/16 [48] CLIP-400M 8$\times$3$\times$4 0.2$\times$12 85.6 97.0
BIKE-L/14 [52] CLIP-400M 16$\times$3$\times$4 0.8$\times$12 88.1 97.9
ActionCLIP-B/16 [34] CLIP-400M 32$\times$3$\times$10 0.6$\times$30 83.8 96.2
X-CLIP-B/16 [34] CLIP-400M 16$\times$3$\times$4 0.6$\times$12 84.7 96.8
S-ViT-B/16 [50] CLIP-400M 16$\times$3$\times$4 0.7$\times$12 84.7 96.8
STAN-B/16 [51] CLIP-400M 16$\times$3$\times$1 0.7$\times$1 84.2 96.5
@@
EVL-B/16 [14] CLIP-400M 8$\times$3$\times$1 0.4$\times$3 82.9 -
ST-Adapter-B/16 [15] CLIP-400M 8$\times$3$\times$1 0.2$\times$3 82.0 95.7
ST-Adapter-B/16 [15] CLIP-400M 32$\times$3$\times$1 0.6$\times$3 82.7 96.2
DUALPATH-B/16 [37] CLIP-400M 32$\times$3$\times$1 0.7$\times$3 85.4 97.1
AIM-B/16 [35] CLIP-400M 8$\times$3$\times$1 0.2$\times$3 83.9 96.3
AIM-B/16 [35] CLIP-400M 32$\times$3$\times$1 0.6$\times$3 84.7 96.7
DiST-B/16 [20] CLIP-400M 8$\times$3$\times$1 0.2$\times$3 83.6 96.3
DiST-B/16 [20] CLIP-400M 32$\times$3$\times$1 0.7$\times$3 85.0 97.0
M2-CLIP-B/16 [18] CLIP-400M 32$\times$3$\times$4 0.8$\times$12 84.1 96.8
Side4Video-B/16 [21] CLIP-400M 32$\times$3$\times$4 0.7$\times$12 84.2 96.5
@@
TDS-CLIP-L/14 CLIP-400M 16$\times$3$\times$4 1.6$\times$12 87.2 97.6
IV-B Comparison with State-of-the-art
IV-B1 Results on Something-Something V1&V2
Tab. I shows the comparison results with other state-of-the-art methods on SSv1 and SSv2. Using both fully fine-tuning and frozen CLIP parameter methods, our approach achieves competitive results on both SSv1 and SSv2. For instance, in the fully fine-tuning approach, UniFormer-V2-L/14 [48] achieves the best classification results with the ViT-L/14 backbone, while our TDS-CLIP demonstrates superior performance with the same backbone (ViT-L/14, 16 frames: 64.1% vs. 60.5% on SSv1; 74.1% vs. 72.1% on SSv2) while using fewer frames. In the frozen CLIP parameter methods, Dist-L/14 [20] uses the same side network structure; on the contrast, TDS-CLIP achieves better results with the same frame and sampling settings (ViT-L/16, 74.1% vs. 72.5% on SSv2), indicating that the proposed adapter enables effective temporal modelling. Additionally, it can be observed that the ViT-L/14 backbone outperforms the ViT-B/16 backbone under the same settings. These experimental results show that the proposed TD-Adapter and SME-Adapter are effective methods for enhancing temporal modeling and facilitating the transfer of knowledge from images to videos in frozen CLIP models.
IV-B2 Results on Kinetics-400
To evaluate the generalization performance of our method, we conduct experiments on the K400 dataset. Compared to SSv1 and SSv2, the K400 emphasizes learning appearance features. In addition, compared with methods such as DiST [20], TDS-CLIP has no independent modeling structure for spatial branches or, instead, simply relies on CLIP’s own powerful spatial modeling capabilities to provide effective spatial features for temporal branches. Therefore, the improvement in K400 dataset is not as obvious as in SSv1 and SSv2, but it still achieves better performance. As shown in Tab. II , our method underperforms BIKE-L/14 [52] in terms of appearance feature learning capability when compared to fully fine-tuned approaches. However, with 8-frame inputs, our performance surpasses that of ActionCLIP-B/16 [34], MViT-B [47], and Video Swin-L [46] with 32-frame inputs, while also maintaining lower computational cost.
When compared to methods that freeze CLIP, our performance with 32-frame inputs is lower than DiST-B/16 [20], AIM-B/16 [35], and DUALPATH-B/16 [37], which possess independent spatial feature learning mechanisms. In contrast, TDS-CLIP focuses more on temporal feature learning and modeling. Notably, we achieve a competitive advantage when compared to methods like DiST-B/16 [20], AIM-B/16 [35], ST-Adapter-B/16 [15], and EVL-B/16 [14] with 8-frame inputs, indicating that TDS-CLIP captures more comprehensive video action information with fewer input frames. Lastly, we also evaluate the performance of TDS-CLIP with an 8-frames and 16-frames ViT-L framework, demonstrating that larger-scale frameworks are effective in improving performance. From the comparison results, we can see that the increase in the number of frames has little gain in performance improvement, which may be the reason for the weak timing correlation of the Kinetics400. Due to memory limitations, we test up to 16 frames of performance in the ViT-L/14, but this is enough to illustrate the advantages. The experimental results demonstrate that TDS-CLIP not only excels in datasets with strong temporal correlations but also performs well in datasets with significant appearance background correlations. With these observations, we can conclude that TDS-CLIP has the dual advantage of spatial modeling and temporal modeling, showcasing robust generalization capabilities.
IV-C Ablation and Analysis
In this section, unless otherwise specifed, we use ViT-B/16 as the backbone and the 8 input frames on Something-Something V1. The experiments used 8 A100 graphics cards and run in the pytorch framework.
IV-C1 Study of component effectiveness
We explore the effectiveness of the proposed adapter and the results are shown in Tab. III. The conventional 3D Conv is used to replace the TD-Adapter for the baseline of comparison. We can observe that the classification accuracy is improved by 1.0% after adding TD-Adapter, which proves that TD-Adapter can effectively improve the temporal modelling capability. Then, we add SME-Adapter, and the accuracy is improved even further, which implies that the SME-Adapter guides the side network to learn the motion features to further enhance the temporal modelling capability of the model, and proves that TD-Adapter and SME-Adapter are complementary to each other in terms of performance improvement.
TABLE III: Study on the location of TD-Adapter layers.
@@ IV-C1 Study of component effectiveness
IV-C2 The impact of local frame sample number
We compare the performance impact of different numbers of local sampling video frames, and the results are shown in Tab.  IV. 1 indicates that the SME-Adapter is not enabled, and serves as the baseline for comparison. The accuracy improves by 0.5% when the number is increased to 3 and by 1.4% when the number is increased to 5. This shows that increasing the number of local sampling frames can effectively improve the recognition performance. We do not test larger quantities because the total number of video frames used for testing is not large and bigger quantities may destroy the sparse sampling efficiency.
TABLE IV: Study on the location of TD-Adapter layers.
@@ IV-C2 The impact of local frame sample number
IV-C3 The impact of fusion rate
The fusion ratio represents the proportion of the proposed method in feature fusion. The test results, as shown in Tab. V, indicate that performance is optimal under equal-ratio fusion, while the performance deteriorates when the output feature ratio of the SME-Adapter is either too high or too low. This shows that the appropriate equal-ratio feature fusion method can well exploit the ability of SME-Adapter to learn and capture motion features.
TABLE V: Study on the location of TD-Adapter layers.
$\alpha$ $\beta$ Top-1(%) Acc.
1.0 - 58.1
- 1.0 47.4
0.5 1.0 58.4
1.0 1.0 59.4
1.0 0.5 58.9
IV-C4 Ablations for TD-Adapter
We explore the impact of different core modules of the TD-Adapter on performance, including temporal convolutional layer enhancement [18] and direct feature difference calculation, as shown in Tab. VI. Using pooling layers achieves a classification accuracy of 59.4%, while temporal convolutional layer enhancement yields a slightly lower accuracy, decreasing by 0.2%. Direct feature difference calculation yields the lowest accuracy, as this approach tends to cause the loss of spatial features. It is worth noting that pooling is a parameter-free operation, thus helping to reduce the number of parameters with essentially the same performance.
Furthermore, we test different kernel sizes for the pooling layers, specifically k=3, 5, and 7, as illustrated in Tab. VII. The classification accuracy decreases as the kernel size increases, indicating that an excessively large kernel size may disrupt the temporal relationships between features, therefore, larger kernel sizes should not be selected.
TABLE VI: Study on the location of TD-Adapter layers.
Method Top-1(%) Acc. GFLOPs
Pool 59.4 -
Conv 59.3 1.7
Direct 57.3 -
TABLE VII: Study on the location of TD-Adapter layers.
@@ IV-C4 Ablations for TD-Adapter
IV-C5 The impact of dimensionality
In the transformer architecture, adjusting the dimensionality can effectively control the model’s memory consumption and impact its performance. Generally, increasing the dimensionality enhances the model’s complexity and memory usage, while also improving the modeling capacity of different layers. We follow the baseline comparison methods to test the impact of various dimensionality settings on performance, as shown in Tab. VIII. The results show that the model achieves the best classification performance at 320 dimensions, while both excessively large or small dimensions constrain performance.
TABLE VIII: Study on the location of TD-Adapter layers.
@@ IV-C5 The impact of dimensionality
IV-C6 Study on the location of TD-Adapter layers
@@
We study the location of the TD-Adapter within the structure, with the results shown in Tab. IX. We experiment with adding the TD-Adapter to the first half of the network (layers 1-6) and the second half (layers 7-12). Replacement with normal 3D convolution when not using TD-Adapter. The table shows that while using more TD-Adapters generally leads to better classification results, adding them to shallower layers provides greater benefits than those obtained from deeper layers. Additionally, the performance is further improved when the SME-Adapter is also used.
TABLE IX: Study on the location of TD-Adapter layers.
Layer 1-6 Layer 7-12 SME-Adapter Top-1 Acc.
58.8%
58.6%
59.1%
59.4%
@@
Figure 5: Visualisation of activation maps for image features. Blue colour indicates larger magnitude of feature activation values and red colour indicates smaller magnitude of feature activation values. White and green boxes indicate focus.
IV-C7 Comparison of modes
In this section, we compare the feature models used by several state-of-the-art methods, as shown in Tab. X. M2-CLIP [18] and DiST [20] model both temporal and spatial branches and use text information. In contrast, Side4video [21] and TDS-CLIP only model temporal branches and do not use text information, which is consistent with the comparison results in Section 4.2 and shows that combining spatial and text information limits certain temporal modeling capabilities but at the same time this is a focus for future research.
TABLE X: Comparison of modes.
Method Mode Text
M2-CLIP [18] Temporal+Sptial
DiST [20] Temporal+Sptial
Side4Video [21] Temporal
TDS-CLIP Temporal
IV-C8 Visualisation Related
@@
As show in Fig. 5, we further visualize the activation maps of features in the side network with and without using the SME-Adapter. It is evident that the SME-Adapter is more sensitive to the motion areas of the target and does not show a bias towards background information (the white box in figure). For example, our SME-Adapter focuses more on the motion region of the hand that interacts with the object, rather than all the targets such as flower pots, water bottles, or mobile phones that appear in the scene (the white green in figure), which is more in line with the human perception. Without the SME-Adapter, the model lacks this motion attention. This indicates that the proposed SME-Adapter facilitates the learning of motion features in videos, enabling the model to focus more accurately on dynamically moving target objects.
Figure 6: Comparison on data efficiency.
IV-C9 SME-Adapter Location Experiment
We test the impact of placing the SME-Adapter at different positions within the network on overall performance. Specifically, we design five modes: (a) Spatial: equipping the SME-Adapter at the input of the CLIP spatial path. (b) Temporal: equipping the SME-Adapter at the input of the side network temporal path. (c) Spatial+Temporal: equipping the SME-Adapter at both the input of the CLIP spatial path and the input of the side network temporal path. (d) Cross Path: connecting the SME-Adapter to the temporal path via a cross-path mode. (e) Additional Path: connecting the SME-Adapter to the temporal path via an additional path mode. The results of these tests are summarized in Tab. XI. We observe that the best performance is achieved when the SME-Adapter is placed solely at the input of the temporal path, whereas equipping it at the CLIP spatial path results in a significant performance degradation. The Cross Path mode is close to the Temporal mode, and Temporal is the best choice for simplicity reasons.
TABLE XI: Study on the location of TD-Adapter layers.
Mode Top-1(%) Acc.
Spatial 53.4
Temporal 60.1
Spatial+Temporal 55.6
Cross Path 59.8
Additional Path 58.7
TABLE XII: Inference Memory Efficiency. w/ indicates “with”.
Component Memory(Gib)
Baseline 8.334
w/ ST-Adapter 23.671
w/ 3D Conv 9.121
w/ TD-Adapter 8.781
w/ SME-Adapter 8.799
TABLE XIII: Configurations for Kinetics-400, Something-Something V1 and Something-Something V2.
dataset Something-Something V1 Something-Something V2 Kinetics-400
backbone ViT-B ViT-L ViT-B ViT-L ViT-B ViT-L
dimension size 320
difference frame number 5
pooling kernel shape 3 $\times$ 1 $\times$ 1
optimizer AdamW, weight decay=0.15, betas=[0.9, 0.999], schedule=Cosine
learning rate 1e-3 1e-4 1e-3 1e-4 1e-3 1e-4
batch size 128 256
epochs 40 30 30 30
warmup epochs 4
seed 1024
training resize RandomResizedCrop
label smoothing 0.1
repeated sampling 2
random augment rand-m7-n4-mstd0.5-inc1
random flip 0.5
test views 3 spatial $\times$ 2 temporal 3 spatial $\times$ 4 temporal
IV-C10 Data Efficiency
Data efficiency refers to fine-tuning a model using only a subset of the training data, thereby maximizing the utility of limited data. Fig. 6 illustrates the impact of different proportions of the training dataset on the performance of TDS-CLIP ViT-B/16. Our model demonstrates superior data efficiency, as evidenced by the B/16 model outperforming EVL ViT-B/16 [14], DiST ViT-B/16 [20], and Side4Video ViT-B/16 [21] by 25.9%, 15.7%, and 2.7% points, respectively, when using only 5% of the training data. Furthermore, our model maintains a competitive advantage across other data efficiency scenarios.
IV-C11 Inference Memory Efficiency
Tab. XII summarizes the memory efficiency in the inference phase. The baseline approach uses only the side network structure without any temporal modeling module, while both temporal and spatial branches are used in ST-Adapter, resulting in a huge memory overhead. From the results, it can be seen that making the proposed method saves a lot of memory.
IV-D Implementation Details
Tab. XIII summarises the fine-tuned configurations for several datasets and models, including Something-Something V1 [13], Something-Something V2 [13] and Kinetics-400 [23]. Something-Something is a large-scale video dataset for action recognition, V1 including about 110K videos (V2 contains more) covering 174 fine-grained action categories. Kinetics-400 is a large-scale YouTube video dataset with about 240K training videos and 20K validation videos covering 400 categories, with each clip lasting about 10 seconds.
All experiments in the paper are implemented in PyTorch [53]. We use CLIP [11] with ViT-B/16, ViT-L/14 for video action recognition. Unless otherwise specified, in general, we use simpler data augmentation techniques than end-to-end fine-tuning.
V Conclusion